OpenStudy (anonymous):

If there is an inverse relationship between x and y, and the constant of variation is forty, what is y when x equals five?

OpenStudy (akashdeepdeb):

So what does inverse mean? That means if x increases y would decrease or x = k*1/y x = 40/y Thus when x = 5 Then 5 = 40/y Now try solving the equation for the y value!! :) Getting me? :D

OpenStudy (anonymous):

@AkashdeepDeb So it 8

OpenStudy (akashdeepdeb):

YES!! :D [Yippe] :)

OpenStudy (jhannybean):

indirect variation: \(\large y = \frac{k}{x}\) , constant of variatoin = \(k = 40\) \[\large y= \frac{40}{x}\] When x=5 \[\large y= \frac{40}{5} = \?\]

OpenStudy (anonymous):

Thanks

OpenStudy (akashdeepdeb):

Being Nice IS NOT flirting.